juan made 2L of lemonade that consists of 40% pure lemon juice. how much water must he add to dilute it to 30% lemon juice
Answer by checkley77(12844)

2*.40=(2+x).30
.80=.60+.30x
.30x=.80-.60
.30x=.20
x=.20/.30
x=2/3 of a liter of water is needed.
Proof:
2*.40=(2+2/3).30
.8=8/3*.30
.8=2.4/3
.8=.8